# Topology of Discrete Quantum Feedback Control

### v2

@@ VI.3.1 Class A
Since a CPTP map always has the modular conjugation symmetry, a CPTP map in a class without $K$ symmetry requires unitary symmetry by which the Bloch matrix can be block-diagonalized, as explained in Sec. V. Here, we explicitly construct such examples of feedback control with unitary symmetry.
Let $\{\hat{M}_{m}\}$ be measurement operators. We assume that the measurement operators commute with an observable $\hat{C}$: $[\hat{M}_{m},\hat{C}]=0\ (\forall m)$. Such measurement operators describe a quantum nondemolition (QND) measurement for the observable $\hat{C}$ in the sense that the expected value of every moment of $\hat{C}$ before and after the measurement stays the same [99], that is,
@@
for arbitrary $\hat{\rho}$ and $r\in\mathbb{N}$. We further assume that the feedback Hamiltonian $\hat{H}_{m}$ commutes with $\hat{C}$, i.e., $[\hat{H}_{m},\hat{C}]=0$ so that $\hat{C}$ is conserved during the feedback control. Then, it follows that the Kraus operator $\hat{K}_{m}=\hat{U}_{m}\hat{M}_{m}$ with $\hat{U}_{m}=\exp[-i\hat{H}_{m}\tau]$ commutes with $\hat{C}$: $[\hat{K}_{m},\hat{C}]=0\ (\forall m)$. Thus, the CPTP map for this feedback control has strong symmetry (see Sec. V.1).
Due to strong symmetry, the CPTP map for the feedback control can be block-diagonalized according to the symmetry eigenvalues for the ket and bra spaces. Let
@@
acts nontrivialily only on the subspace $\mathcal{H}_{c}\otimes\mathcal{H}_{c^{\prime}}$.
The BL symmetry classes are examined for each block-diagonalized matrix ~$\tilde{\mathcal{E}}_{c,c^{\prime}}$. A crucial observation here is that each block does not necessarily have the modular conjugation symmetry since
@@
In particular, ~$\tilde{\mathcal{E}}_{c,c^{\prime}}$ with $c\neq c^{\prime}$ does not possess the modular conjugation symmetry, but it is related to ~$\tilde{\mathcal{E}}_{c^{\prime},c}$ by this symmetry. Thus, it belongs to class A if it does not have any other symmetry.
@@ VI.3.2 Class AI
The other classes (psH, AI, and AII) without $K$ symmetry include $C$ or $Q$ symmetry. These symmetries are realized with feedback protocols with two-point measurements similar to those in Sec. VII.2. Specifically, we consider a spin-$1/2$ particle with two sublattice degrees of freedom, where a quantum state at site $\vec{j}$ with sublattice $a=A,B$ and spin $\sigma=\uparrow,\downarrow$ is denoted by |$\ket{\vec{j},a,\sigma}$. We perform a projective position measurement on this particle, and assume that this measurement does not disturb the spin state of the particle. Thus, a measurement outcome is specified only by the site index $\vec{j}$ and the sublattice index $a$, and the measurement operator is given by
@@
where $\hat{K}_{m_{1},m_{2}}=\hat{M}_{m_{2}}\hat{U}_{m_{1}}\hat{M}_{m_{1}}$ is the Kraus operator, $\hat{M}_{m}$ is the measurment operator (97) corresponding to a measurement outcome $m=(\vec{j},a)$, and $\hat{U}_{m}$ is a unitary operator for a feedback operation. We assume that $[\hat{U}_{m},\hat{S}^{z}]=0$. As in the case of a single measurement in Sec. VI.3.1, the CPTP map can be decomposed as
@@
This condition is interpreted as reciprocity in the feedback process. Since the truncated Bloch matrix $X_{\mathrm{trunc}}^{(\uparrow\downarrow)}(\vec{k})$ does not have $K$ symmetry, it belongs to the symmetry class AI.
@@ VI.3.3 Class AII
The symmetry class AII involves $C$ symmetry with $\eta_{C}=-1$. Provided that
@@ VI.3.4 Class psH
The psH (112) is understood as a consequence of the modified reciprocity expressed by Eqs. (115a)-(115c).
VI.4 Class AII, AII $+$ psH+, and AII $+$ psH-: feedback with $K$ symmetry and $\eta_{K}=-1$
@@
We note that these measurement operators make the spin degrees of freedom unaltered. We assume that feedback Hamiltonians conserve the magnetization. Then, a CPTP map for feedback control with a two-point measurement is given by the form (99), where the measurement outcomes are labeled by $m_{n}=(\vec{j}_{n},a_{n},\ell_{n})\ (n=1,2)$. The truncated Bloch matrix is obtained in a manner similar to that in Sec. VI.3.
@@
VII Topological Maxwell’s demons
In this section, builiding on the symmetry classification of quantum channels developed in Sec. V, we present various models of topological feedback control. In Sec. VII.1-VII.4, we focus on quantum feedback control of a particle in a 1D lattice and construct topological quantum channels protected by their point-gap topology and symmetry. According to the general topological classification of non-Hermitian operators in Ref. [34], a non-Hermitian operator in a 1D system can host nontrivial point-gap topology in classes A, AI, AI $+$ psH-, AII, and AII. In Sec. VII.1, we consider topological feedback control in class AI, where we can combine particle-hole symmetry with the modular conjugation symmmetry to construct a unitary symmetry. In Sec. VII.2, we construct a model of topological feedback control in class AI $+$ psH- and demonstrate that it shows helical spin transport protected by $C$ symmetry. In Sec. VII.3, we consider topological feedback control in class A, where strong unitary symmetry enables us to block-diagonalize a CPTP map. In Sec. VII.4, we impose strong unitary symmetry and $C$ symmetry on a CPTP map to construct topological feedback control in class AII, which exhibits the $\mathbb{Z}_{2}$ skin effect. Finally, in Sec. VII.5, we show an example of quantum feedback control with nontrivial line-gap topology in two dimensions.
@@ VII.1 Su-Schrieffer-Heeger demon
First, we present a model of topological feedback control with $K$ symmetry with $\eta_{K}=+1$ that is distinct from the modular conjugation symmetry. This model provides a concrete example for the discussion presented in Sec. VI.1. We consider a 1D lattice with two sublattices. Each unit cell is labeled by an index $j$ and each sublattice is labeled by $a=A,B$. We first perform projective position measurement which completely resolves the sublattice structure. A measurement outcome is specified by $m=(j,a)$, and the corresponding measurement operator is given by $\hat{M}_{m}=\ket{j,a}\bra{j,a}$ |. Then, we perform feedback control with Hamiltonian $\hat{H}_{a}\ (a=A,B)$ that depends only on the measurement outcome of the sublattice index. The feedback Hamiltonians are given by the Su-Schrieffer-Heeger (SSH) model [100]:
@@
which follow from Eqs. (124a), (124b), and (124c), respectively. Among the symmetries (126a)-(126c), only particle-hole symmetry (126b) leads to the symmetry of a quantum channel for this feedback control. Since the measurement operator satisfies
@@
Figure 12: Schematic illustration of the Su-Schrieffer-Heeger (SSH) demon. We first perform projective measurement of the position and the sublattice. (a) If a particle is detected in sublattice $A$, we perform feedback control governed by the Hamiltonian $H_{A}$. (b) If a particle is detected in sublattice $B$, we perform feedback control governed by the Hamiltonian $H_{B}$.
As mentioned in Sec. V, the CPTP map has the modular conjugation symmetry (62), which gives another $K$ symmetry different from (128). Combining the two $K$ symmetries, we obtain the unitary symmetry of the CPTP map as
$(\Sigma_{z}\otimes\Sigma_{z}^{*})\tilde{S}\sum_{m}(\hat{K}_{m}\otimes\hat{K}_{m}^{*})\tilde{S}^{-1}(\Sigma_{z}\otimes\Sigma_{z}^{*})=\sum_{m}(\hat{K}_{m}\otimes\hat{K}_{m}^{*}),$ (129)
where $\tilde{S}$ is the swap operator defined by Eq. (72). Hence, the CPTP map can be block-diagonalized according to the symmetry eigenvalues of the unitary symmetry (129). The eigenvalues of $(\hat{\Sigma}_{z}\otimes\hat{\Sigma}_{z})\tilde{S}$ are given by ±$\pm 1$. The eigenvectors for the eigenvalue $+1$ are given by
@@
In each eigenspace of the operator $(\hat{\Sigma}_{z}\otimes\hat{\Sigma}_{z})\tilde{S}$, the modular conjugation symmetry (62) and the particle-hole symmetry (128) are equivalent to each other since $(\hat{\Sigma}_{z}\otimes\hat{\Sigma}_{z})=\pm\tilde{S}$. Thus, the topological characterization of a CPTP map in those eigenspaces reduces to that of a CPTP map with a single $K$ symmetry. Thus, each symmetry sector of the CPTP map belongs to class AI.
Figure 13 shows the eigenspectrum of a CPTP map for the feedback control with the SSH model under the PBC and that under the OBC. The eigenspectrum under the PBC forms two loops in the complex plane, each of which has the winding number $-1$ for a point gap inside the loop. This nontrivial point-gap topology makes the eigenspectrum under the OBC drastically different from that in the PBC, leading to the non-Hermitian skin effect induced by the SSH demon.
@@
VII.2 Helical Maxwell’s demon
@@
where $J\in\mathbb{R}$ is the hopping amplitude, $V>0$ is the height of the potential, and $h_{x}$ denotes the strength of a transverse magnetic field which mixes the spin components. The feedback control is described by a unitary operator $\hat{U}_{j_{1},\sigma_{1}}=\exp[-i\hat{H}_{j_{1},\sigma_{1}}\tau]$. After the feedback control, we perform the second projective measurement with projection operators $\hat{P}_{j,\sigma}$. Let $m_{2}=(j_{2},\sigma_{2})$ be an outcome of this second measurement. If $\sigma_{2}=\sigma_{1}$, we do nothing. If $\sigma_{2}\neq\sigma_{1}$, we flip the spin of the particle by applying a $\pi$ pulse of a magnetic field. Through this feedback control, a spin-up particle is transported to the right, while a spin-down particle is transported to the left, which is analogous to helical edge states in the quantum spin Hall effect [12, 13, 14, 15, 16]
Figure 14: Schematic illustration of the helical Maxwell’s demon for a spin-$1/2$ particle on a 1D lattice. We first perform a projective measurement of the position and spin of the particle. If we obtain an outcome $m_{1}=(j_{1},\uparrow)$ ($m_{1}=(j_{1},\downarrow)$), we raise the potential at site $j_{1}-1$ ($j+1$) and let the particle undergo unitary time evolution under the feedback Hamiltonian given in Eq. (132) (Eq. (133)). After that, we perform the second projective measurement of the position and spin. If the measurement outcome of spin is different from that of the first measurement, we apply a $\pi$ pulse of a magnetic field to flip the spin with unit probability. Thus, helical spin transport is achieved.
@@
Note that we can transfer the off-diagonal components in Eq. (80) to the diagonal components by the additional feedback control after the second measurement. The truncated Bloch matrix has the symmetry
@@
where $\xi_{\mathrm{PG}}$ denotes the location of a point gap and we use Eqs. (140a) and (140b) in deriving the second equality. In Fig. 15(a), we show the eigenspectra of the CPTP map under the PBC and the OBC. Each eigenvalue is doubly degenerate due to the Kramers degeneracy ensured by the $C$ symmetry with $\eta_{C}=-1$ [34]. The loop structure of the eigenspectrum under the PBC leads to the winding number $w_{\mathrm{spin}}(\xi_{\mathrm{PG}})=-1$ for a point gap $\xi_{\mathrm{PG}}$ inside the loop. Reflecting the nontrivial point-gap topology, the eigenspectrum under the OBC is drastically changed from that under the PBC, and the right eigenmodes of the spin-up (spin-down) component are localized near the right (left) edge of the system. This is a manifestation of the symmetry-protected non-Hermitian skin effect [72], where the Kramers doublets are localized near the opposite edges. Physically, the symmetry-protected non-Hermitian skin effect is a consequence of the helical spin transport due to feedback control. Thus, the helical Maxwell’s demon considered here achieves robust feedback-assisted spin transport protected by nontrivial topology of a quantum channel.
The point-gap topology of the helical Maxwell’s demon is protected by the $C$ symmetry. To see this, we show the eigenspectrum of another CPTP map
@@
where the difference from Eq. (134) is the absense of the feedback after the second measurement. The truncated Bloch matrix of this CPTP map is given by
@@
where the hybridization between spin components implies the absense of the $C$ symmetry with $\eta_{C}=-1$. The eigenspectra of the CPTP map $\mathcal{E}^{\prime}$ are shown in Fig. 15(b). Due to the lack of the $C$ symmetry, the Kramers degeneracy of the eigenvalues no longer holds. In fact, the two degenerate loops of eigenvalues for $h_{x}=0$ are hybridized due to the real-complex transition near the steady state. As a consequence of spin mixing, the non-Hermitian skin effect disappears under the OBC and hence the topology becomes trivial. Thus, the symmetry protection of the topological feedback control is demonstrated.
Figure 15: (a) Eigenspectrum of a quantum channel for the helical Maxwell’s demon with additional feedback performed after the second measurement. (b) Eigenspectrum of a quantum channel for the helical Maxwell’s demon with no further feedback performed after the second measurement. An enlarged eigenspectrum near the steady state is also shown. Blue curves and the blue dot at the origin constitute the eigenspectrum under the PBC, and orange (green) dots form the eigenspectrum under the OBC with $L=30$ ($L=20$). The hopping amplitude is set to $J=1$ and the transverse magnetic field is set to $h_{x}=0.2$. The duration of feedback is $\tau=1$.

### v3

@@ VI.3.1 Class A
Since a CPTP map always has modular conjugation symmetry, a CPTP map in a class without $K$ symmetry requires unitary symmetry by which the Bloch matrix can be block diagonalized, as explained in Sec. V. Here, we explicitly construct such examples of feedback control with unitary symmetry.
Let $\{\hat{M}_{m}\}$ be measurement operators. We assume that the measurement operators commute with an observable $\hat{C}$: $[\hat{M}_{m},\hat{C}]=0\ (\forall m)$. Such measurement operators describe a quantum nondemolition (QND) measurement for the observable $\hat{C}$ in the sense that all the moments of $\hat{C}$ before and after the measurement stay the same [99], that is,
@@
for arbitrary $\hat{\rho}$ and $r\in\mathbb{N}$. We further assume that the feedback Hamiltonian $\hat{H}_{m}$ commutes with $\hat{C}$, i.e., $[\hat{H}_{m},\hat{C}]=0$ so that $\hat{C}$ is conserved during feedback control. Then, it follows that the Kraus operator $\hat{K}_{m}=\hat{U}_{m}\hat{M}_{m}$ with $\hat{U}_{m}=\exp[-i\hat{H}_{m}\tau]$ commutes with $\hat{C}$: $[\hat{K}_{m},\hat{C}]=0\ (\forall m)$. Thus, the CPTP map for this feedback control has strong symmetry (see Sec. V.1).
Because of strong symmetry, the CPTP map for feedback control can be block diagonalized according to the symmetry eigenvalues for the ket and bra spaces. Let
@@
acts nontrivially only on the subspace $\mathcal{H}_{c}\otimes\mathcal{H}_{c^{\prime}}$.
The BL symmetry classes are examined for each block diagonalized matrix ~$\tilde{\mathcal{E}}_{c,c^{\prime}}$. A crucial observation here is that every block does not necessarily have the modular conjugation symmetry since
@@
In particular, ~$\tilde{\mathcal{E}}_{c,c^{\prime}}$ with $c\neq c^{\prime}$ does not possess modular conjugation symmetry, but it is related to ~$\tilde{\mathcal{E}}_{c^{\prime},c}$ by this symmetry. Thus, it belongs to class A if it does not have any other symmetry.
@@ VI.3.2 Class AI
The other classes (psH, AI, and AII) without $K$ symmetry include $C$ or $Q$ symmetry. These symmetries are realized with feedback protocols with two-point measurements similar to those in Sec. VII.2. Specifically, we consider a spin-$1/2$ particle with two sublattice degrees of freedom, where a quantum state at site $\vec{j}$ with sublattice $a=A,B$ and spin $\sigma=\uparrow,\downarrow$ is denoted by |$\ket{\vec{j},a,\sigma}$. We perform a projective position measurement on this particle and assume that this measurement does not disturb the spin state of the particle. Thus, a measurement outcome is specified only by the site index $\vec{j}$ and the sublattice index $a$, and the measurement operator is given by
@@
where $\hat{K}_{m_{1},m_{2}}=\hat{M}_{m_{2}}\hat{U}_{m_{1}}\hat{M}_{m_{1}}$ is the Kraus operator, $\hat{M}_{m}$ is the measurement operator (97) corresponding to a measurement outcome $m=(\vec{j},a)$, and $\hat{U}_{m}$ is a unitary operator for feedback operation. We assume that $[\hat{U}_{m},\hat{S}^{z}]=0$. As in the case of a single measurement in Sec. VI.3.1, the CPTP map can be decomposed as
@@
This condition is interpreted as reciprocity in the feedback process. Since the truncated Bloch matrix $X_{\mathrm{trunc}}^{(\uparrow\downarrow)}(\vec{k})$ does not have $K$ symmetry, it belongs to symmetry class AI.
@@ VI.3.3 Class AII
Symmetry class AII involves $C$ symmetry with $\eta_{C}=-1$. Provided that
@@ VI.3.4 Class psH
The psH in Eq. (112) is understood as a consequence of the modified reciprocity expressed by Eqs. (115a)-(115c).
VI.4 Classes AII, AII $+$ psH+, and AII $+$ psH-: Feedback with $K$ symmetry and $\eta_{K}=-1$
@@
We note that these measurement operators do not alter the spin degrees of freedom. We assume that feedback Hamiltonians conserve the magnetization. Then, a CPTP map for feedback control with a two-point measurement is given by the form (99), where the measurement outcomes are labeled by $m_{n}=(\vec{j}_{n},a_{n},\ell_{n})\ (n=1,2)$. The truncated Bloch matrix is obtained in a manner similar to that in Sec. VI.3.
@@
VII topological Maxwell demons
In this section, building on the symmetry classification of quantum channels developed in Sec. V, we present various models of topological feedback control. In Secs. VII.1-VII.4, we focus on quantum feedback control of a particle in a 1D lattice and construct topological quantum channels protected by their point-gap topology and symmetry. According to the general topological classification of non-Hermitian operators in Ref. [34], a non-Hermitian operator in a 1D system can host nontrivial point-gap topology in classes A, AI, AI $+$ psH-, AII, and AII. In Sec. VII.1, we consider topological feedback control in class AI, where we can combine particle-hole symmetry with modular conjugation symmetry to construct a unitary symmetry. In Sec. VII.2, we construct a model of topological feedback control in class AI $+$ psH- and demonstrate that it shows helical spin transport protected by $C$ symmetry. In Sec. VII.3, we consider topological feedback control in class A, where strong unitary symmetry enables us to block diagonalize a CPTP map. In Sec. VII.4, we impose strong unitary symmetry and $C$ symmetry on a CPTP map to construct topological feedback control in class AII, which exhibits the $\mathbb{Z}_{2}$ skin effect. Finally, in Sec. VII.5, we show an example of quantum feedback control with nontrivial line-gap topology in two dimensions.
@@ VII.1 Su-Schrieffer-Heeger demon
First, we present a model of topological feedback control with $K$ symmetry with $\eta_{K}=+1$ that is distinct from the modular conjugation symmetry. This model provides a concrete example for the discussion presented in Sec. VI.1. We consider a 1D lattice with two sublattices. Each unit cell is labeled by an index $j$ and each sublattice is labeled by $a=A,B$. We first perform a projective position measurement, which completely resolves the sublattice structure. A measurement outcome is specified by $m=(j,a)$, and the corresponding measurement operator is given by $\hat{M}_{m}=\ket{j,a}\bra{j,a}$ |. Next, we perform feedback control with Hamiltonian $\hat{H}_{a}\ (a=A,B)$, which depends only on the measurement outcome of the sublattice index. The feedback Hamiltonians are given by the Su-Schrieffer-Heeger (SSH) model [100]:
@@
which follow from Eqs. (124a)-(124c), respectively. Among the symmetries (126a)-(126c), only particle-hole symmetry (126b) leads to the symmetry of a quantum channel for this feedback control. Since the measurement operator satisfies
@@
Figure 12: Schematic illustration of SSH demon. We first perform a projective measurement of the position and the sublattice. In panel (a), we show that if a particle is detected in sublattice $A$, we perform feedback control governed by the Hamiltonian $H_{A}$. Panel (b) shows that if a particle is detected in sublattice $B$, we perform feedback control governed by the Hamiltonian $H_{B}$.
As mentioned in Sec. V, the CPTP map has modular conjugation symmetry (62), which gives another $K$ symmetry different from (128). Combining the two $K$ symmetries, we obtain the unitary symmetry of the CPTP map as
$(\hat{\Sigma}_{z}\otimes\hat{\Sigma}_{z}^{*})\tilde{S}\sum_{m}(\hat{K}_{m}\otimes\hat{K}_{m}^{*})\tilde{S}^{-1}(\hat{\Sigma}_{z}\otimes\hat{\Sigma}_{z}^{*})=\sum_{m}(\hat{K}_{m}\otimes\hat{K}_{m}^{*}),$ (129)
where $\tilde{S}$ is the swap operator defined by Eq. (72). Hence, the CPTP map can be block diagonalized according to the symmetry eigenvalues of the unitary symmetry (129). The eigenvalues of $(\hat{\Sigma}_{z}\otimes\hat{\Sigma}_{z}^{*})\tilde{S}$ are given by ±$\pm 1$. The eigenvectors for the eigenvalue $+1$ are given by
@@
In each eigenspace of the operator $(\hat{\Sigma}_{z}\otimes\hat{\Sigma}_{z}^{*})\tilde{S}$, modular conjugation symmetry (62) and particle-hole symmetry (128) are equivalent to each other since $(\hat{\Sigma}_{z}\otimes\hat{\Sigma}_{z}^{*})=\pm\tilde{S}$. Thus, the topological characterization of a CPTP map in those eigenspaces reduces to that of a CPTP map with a single $K$ symmetry. Thus, each symmetry sector of the CPTP map belongs to class AI.
Figure 13 shows the eigenspectrum of a CPTP map for feedback control with the SSH model under the PBC as well as that under the OBC. The eigenspectrum under the PBC forms two loops in the complex plane, each of which has winding number $-1$ for a point gap inside the loop. This nontrivial point-gap topology makes the eigenspectrum under the OBC drastically different from that in the PBC, leading to the non-Hermitian skin effect induced by the SSH demon.
@@
VII.2 Helical Maxwell demon
@@
where $J\in\mathbb{R}$ is the hopping amplitude, $V>0$ is the height of the potential, and $h_{x}$ denotes the strength of a transverse magnetic field that mixes the spin components. Feedback control is described by a unitary operator $\hat{U}_{j_{1},\sigma_{1}}=\exp[-i\hat{H}_{j_{1},\sigma_{1}}\tau]$. After feedback control, we perform the second projective measurement with projection operators $\hat{P}_{j,\sigma}$. Let $m_{2}=(j_{2},\sigma_{2})$ be an outcome of this second measurement. If $\sigma_{2}=\sigma_{1}$, we do nothing. If $\sigma_{2}\neq\sigma_{1}$, we flip the spin of the particle by applying a $\pi$ pulse of the magnetic field. Through this feedback control, a spin-up particle is transported to the right while a spin-down particle is transported to the left, which is analogous to helical edge states in the quantum spin Hall effect [12, 13, 14, 15, 16].
Figure 14: Schematic illustration of the helical Maxwell demon for a spin-$1/2$ particle on a 1D lattice. We first perform a projective measurement of the position and spin of the particle. If we obtain an outcome $m_{1}=(j_{1},\uparrow)$ ($m_{1}=(j_{1},\downarrow)$), we raise the potential at site $j_{1}-1$ ($j+1$) and let the particle undergo unitary time evolution under the feedback Hamiltonian given in Eq. (132) [Eq. (133)]. Next, we perform the second projective measurement of the position and spin. If the measurement outcome of spin is different from that of the first measurement, we apply a $\pi$ pulse of a magnetic field to flip the spin with unit probability. Thus, helical spin transport is achieved.
@@
Note that we can transfer the off-diagonal components in Eq. (80) to the diagonal components by an additional feedback control after the second measurement. The truncated Bloch matrix has symmetry
@@
where $\xi_{\mathrm{PG}}$ denotes the location of a point gap and we use Eqs. (140a) and (140b) in deriving the second equality. In Fig. 15(a), we show the eigenspectra of the CPTP map under the PBC and the OBC. Each eigenvalue is doubly degenerate due to the Kramers degeneracy ensured by $C$ symmetry with $\eta_{C}=-1$ [34]. The loop structure of the eigenspectrum under the PBC leads to the winding number $w_{\mathrm{spin}}(\xi_{\mathrm{PG}})=-1$ for a point gap $\xi_{\mathrm{PG}}$ inside the loop. Reflecting the nontrivial point-gap topology, the eigenspectrum under the OBC is drastically changed from that under the PBC, and the right eigenmodes of the spin-up (spin-down) component are localized near the right (left) edge of the system. This phenomenon is a manifestation of the symmetry-protected non-Hermitian skin effect [72], where the Kramers doublets are localized near the opposite edges. Physically, the symmetry-protected non-Hermitian skin effect is a consequence of the helical spin transport due to feedback control. Thus, the helical Maxwell demon considered here achieves robust feedback-assisted spin transport protected by nontrivial topology of a quantum channel.
The point-gap topology of the helical Maxwell demon is protected by $C$ symmetry. To see this property, we show the eigenspectrum of another CPTP map,
@@
where the difference from Eq. (134) is the absence of feedback after the second measurement. The truncated Bloch matrix of this CPTP map is given by
@@
where the hybridization between spin components implies the absence of $C$ symmetry with $\eta_{C}=-1$. The eigenspectra of the CPTP map $\mathcal{E}^{\prime}$ are shown in Fig. 15(b). Because of the lack of $C$ symmetry, the Kramers degeneracy of the eigenvalues no longer holds. In fact, the two degenerate loops of eigenvalues for $h_{x}=0$ are hybridized due to the real-complex transition near the steady state. As a consequence of spin mixing, the non-Hermitian skin effect disappears under the OBC and hence the topology becomes trivial. Thus, the symmetry protection of topological feedback control is demonstrated.
Figure 15: (a) Eigenspectrum of a quantum channel for the helical Maxwell demon with additional feedback performed after the second measurement. (b) Eigenspectrum of a quantum channel for the helical Maxwell demon with no further feedback performed after the second measurement. An enlarged eigenspectrum near the steady state is also shown. Blue curves and the blue dot at the origin constitute the eigenspectrum under the PBC, and orange (green) dots form the eigenspectrum under the OBC with $L=30$ ($L=20$). The hopping amplitude is set to $J=1$, and the transverse magnetic field is set to $h_{x}=0.2$. The duration of feedback is $\tau=1$.